# Ultra-high-energy event KM3-230213A constraints on Lorentz Invariance Violation in neutrino sector

### v1

@@ Abstract
We discuss the constraints on superluminal neutrino Lorentz Invariance Violation (LIV) parameters from the observation of the ultra-high-energy event KM3-230213A by KM3NeT collaboration in cases of linear $n=1$ and quadratic $n=2$ LIV scenarios. Assuming extragalactic origin of the event, we obtain the constraints on LIV mass scale $\Lambda_{n=1}=5.4\times 10^{30}\,\mbox{GeV}$ and $\Lambda_{n=2}=3.5\times 10^{19}\,\mbox{GeV}$ from the absence of neutrino splitting.
@@ Introduction
Recently KM3NeT collaboration reported [1] the event called KM3-230213A that has been identified with astrophysical neutrino of energy $220^{+570}_{-100}$ PeV. This energy is almost two orders of magnitude larger than the previous most energetic neutrino event detected at IceCube. The origin of the neutrino event KM3-230213A is an intriguing question: galactic [2], extragalactic [3] and cosmogenic [4] possible origins are discussed. Such ultra high energy of the event allows us to test some fundamental principles such as Lorentz Invariance at an unprecedented level. Small departures of LI appear in some models of quantum gravity, see [5] for a review.
@@
where $n=1$ and $n=2$ respond to linear and quadratic LIV correspondingly; $\Lambda$ is the energy scale at which LI is broken. $s_{n}$ is the sign which strongly influences the LIV phenomenology: $s_{n}=+1\,(-1)$ called superluminal (subluminal) LIV which lead to superluminal (subluminal) velocity of a given particle. Besides, LIV leads to the modification of thresholds and cross-sections of particle processes; superluminal LIV in general leads to several decay channels of a given particle absent in LI case; subluminal LIV, in contrast, suppresses some interactions present in LI.
In this letter we show that the detection of the neutrino event leads to neutrino (meta)stability (or long lifetime) which in turn leads to the absence of concrete decay channels. Consequently, we limit ourselves to consider superluminal neutrino. The corresponding decay channels are neutrino splitting $\nu\to\nu\nu\bar{\nu}$, discussed in [6, 7] and neutrino pair production $\nu\to\nu e^{+}e^{-}$. Before the numerical consideration of the processes, let us provide the best current bound [^1] Originally $c_{6}>-5.2\cdot 10^{-35}\ \mbox{GeV}^{-2}$ in terms of the Data Tables [8]. on superluminal LIV from IceCube [9],
@@
In order to consider cross-sections of neutrino-involved processes one should start from effective field theory (EFT) Lagrangian. The minimal LIV addition to the Standard Model (SM) Lagrangian, preserving $SU(2)_{L}$ gauge symmetry reads,
@@
where $l_{L}^{\alpha}=(\nu_{\alpha},l_{\alpha})_{L}$ is the left lepton doublet, $\alpha$ counts the lepton generations, $D_{0}$ is the time part of the SM covariant derivative. Note that $n=1$ scenario is CPT-odd while $n=2$ is CPT-even. Considering only quadratic over neutrino part of (3), one obtains the dispersion relation,
@@
In order to take a numerical value of the neutrino width associated with neutrino splitting we follow the analytical result of Carmona et al [7] obtained in the collinear approximation neglecting neutrino masses. The width of the neutrino splitting does not depend on flavour and reads [7],
@@
Here $c^{(\nu)}\approx 0.024$, $G_{F}$ is the Fermi constant. More detailed characteristic is the neutrino energy losses, see [7]. We leave this approach for future more detailed analysis. Solving eq. (6) with respect to $\Lambda$, one obtains
@@
where $L=\Gamma_{\nu\to\nu\nu\bar{\nu}}^{-1}$ is the mean free path of neutrino associated to the splitting. We take it as the distance to the possible source.
Further we use the median value of KM3-230213A energy, $E=220$ PeV, and consider three scenario of neutrino origin: galactic, extragalactic and cosmogenic ones. In the galactic scenario, we take $L=10$ kpc, so eq. (7) reads,
$\Lambda=\begin{cases}5.4\times 10^{29}\,\mbox{GeV},&n=1,\\ 1.1\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (8)
@@
$\Lambda=\begin{cases}5.4\times 10^{30}\,\mbox{GeV},&n=1,\\ 3.5\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (9)
@@
$\Lambda=\begin{cases}1.2\times 10^{31}\,\mbox{GeV},&n=1,\\ 5.1\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (10)
Splitting cannot produce significant excess at lower energy in a spectrum decreasing with energy like $\sim E^{-2}$. In fact, splitting of neutrino of energy $E$ into 3 particles of energy $E/3$ add 3 particles to the spectrum which consists of $(E/(E/3))^{2}$ = 9 particles.
@@
For aforementioned values of $\Lambda$ and $E$ one cannot neglect the electron mass values, so one should estimate the threshold effect, which is in general complicated for non-equal 3 particles in the final state. Consider collinear kinematics, neutrino of energy and x-projection of momentum $(k+\frac{k^{3}}{2\Lambda^{2}},k)$, to neutrino, electron a positron each of them takes $k/3$ momentum. Finally, the threshold of the effective decay reads,
$\Lambda_{thr}=\frac{E^{2}}{\sqrt{6}m_{e}}=4.0\times 10^{19}\,\mbox{GeV},\qquad n=2.$ (11)
For larger $\Lambda$ the process does not occur, for lower $\Lambda$ the calculation are similar to the neutrino splitting. Carmona et al [7] calculated the corresponding width in case of zero electron mass; the width is that of eq. (7) multiplied by $0.13$. Finally, for our values of parameters $\nu\to\nu e^{+}e^{-}$ is subleading with respect to neutrino splitting.
@@
Consider the opposite process, the production of muon neutrino. In the majority of models neutrino is produced very fast in the process of pion decay, $\pi^{+}\to\mu^{+}\nu_{\mu}$. If the process is kinematically forbidden, no neutrino are produced. Two-particle decay threshold rule reads,
@@
From that we obtain we constraint,
$\Lambda>5.1\times 10^{18}\,\mbox{GeV},\qquad n=2.$ (13)
@@ Conclusion
We estimated constraints on LIV in the neitrino sector from the event KM3-230213A. It turns out that the best constraints came from the absence of the neutrino splitting. We considered separately the cases for neutrino of galactic, extragalactic and cosmogenic origin. Since it is concluded in [2] that the galactic origin is unlikely, we may take the constraint (9) as the conservative one, precisely
$\Lambda=\begin{cases}5.4\times 10^{30}\,\mbox{GeV},&n=1,\\ 3.5\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (14)
In order to obtain more precise statistical significant constraints, the models of galactic, extragalactic or cosmogenic sources are required as well as an accurate calculation of neutrino energy losses.

### v3

@@ Abstract
We discuss the constraints on superluminal neutrino Lorentz Invariance Violation (LIV) parameters from the observation of the ultra-high-energy event KM3-230213A by the KM3NeT collaboration in cases of linear $n=1$ and quadratic $n=2$ LIV scenarios. Assuming extragalactic origin of the event, we obtain the constraints on LIV mass scale $\Lambda_{n=1}=1.1\times 10^{30}\,\mbox{GeV}$ and $\Lambda_{n=2}=1.1\times 10^{19}\,\mbox{GeV}$ from the absence of neutrino splitting.
preprint: INR-TH-2025-001
@@ Introduction
Recently, KM3NeT collaboration reported [1] the event called KM3-230213A that has been identified with astrophysical neutrino of energy $220^{+570}_{-100}$ PeV. This energy is almost two orders of magnitude larger than the previous most energetic neutrino event detected at IceCube. The origin of the neutrino event KM3-230213A is an intriguing question: galactic [2], extragalactic [3] and cosmogenic [4] possible origins are discussed. Such ultra high energy of the event allows us to test some fundamental principles such as Lorentz Invariance at an unprecedented level. Small departures of LI appear in some models of quantum gravity; see [5] for a review.
@@
where $n=1$ and $n=2$ respond to linear and quadratic LIV correspondingly; $\Lambda$ is the energy scale at which LI is broken. $s_{n}$ is the sign which strongly influences the LIV phenomenology: $s_{n}=+1\,(-1)$ called superluminal (subluminal) LIV which lead to superluminal (subluminal) velocity of a given particle. Besides, LIV leads to the modification of thresholds and cross-sections of particle processes; superluminal LIV in general leads to several decay channels of a given particle absent in the LI case; subluminal LIV, in contrast, suppresses some interactions present in LI.
In this letter we show that the detection of the neutrino event leads to neutrino (meta)stability (or long lifetime) which in turn leads to the absence of concrete decay channels. Consequently, we limit ourselves to consider superluminal neutrino. The corresponding decay channels are neutrino splitting $\nu\to\nu\nu\bar{\nu}$, discussed in [6, 7, 8] and neutrino pair production $\nu\to\nu e^{+}e^{-}$ [9, 10, 6]. Before the numerical consideration of the processes, let us provide the best current bound [^1] Originally $c_{6}>-5.2\cdot 10^{-35}\ \mbox{GeV}^{-2}$ in terms of the Data Tables [11]. on superluminal LIV from IceCube [6],
@@
In order to consider cross-sections of neutrino-involved processes, one should start from the effective field theory (EFT) Lagrangian. The minimal LIV addition to the Standard Model (SM) Lagrangian, preserving $SU(2)_{L}$ gauge symmetry reads,
@@
where $l_{L}^{\alpha}=(\nu_{\alpha},l_{\alpha})_{L}$ is the left lepton doublet, $\alpha$ counts the lepton generations, $D_{0}$ is the time part of the SM covariant derivative. Note that $n=1$ scenario is CPT-odd while $n=2$ is CPT-even. Considering only the quadratic over neutrino part of (3), one obtains the dispersion relation,
@@
In order to take a numerical value of the neutrino width associated with neutrino splitting, we follow the analytical result of Carmona et al [8] obtained in the collinear approximation neglecting neutrino masses. The width of the neutrino splitting does not depend on flavour and reads [8],
@@
Here $c^{(\nu)}\approx 0.024$, $G_{F}$ is the Fermi constant. More detailed characteristics are the neutrino energy losses, see [8]. We leave this approach for a future more detailed analysis. Solving eq. (6) with respect to $\Lambda$, one obtains
@@
where $L=\Gamma_{\nu\to\nu\nu\bar{\nu}}^{-1}$ is the mean free path of the neutrino associated with the splitting. We take it as the distance to the possible source.
Further we use the median value of KM3-230213A energy with $68\%$ CL errorbars, $E=220^{+570}_{-100}$ PeV, and consider three scenarios of neutrino origin: galactic, extragalactic and cosmogenic ones. In the galactic scenario, we take $L=10$ kpc, so eq. (7) reads,
$\Lambda=\begin{cases}5.3^{+156.2}_{-4.2}\times 10^{29}\,\mbox{GeV},&n=1,\\ 1.1_{-0.7}^{+10.2}\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (8)
@@
$\Lambda=\begin{cases}5.3^{+156.2}_{-4.2}\times 10^{30}\,\mbox{GeV},&n=1,\\ 3.4^{+32.3}_{-2.3}\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (9)
@@
$\Lambda=\begin{cases}1.2^{+33.6}_{-1.0}\times 10^{31}\,\mbox{GeV},&n=1,\\ 5.0^{+47.4}_{-3.3}\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (10)
Splitting cannot produce significant excess at lower energy in a spectrum decreasing with energy like $\sim E^{-2}$. In fact, splitting of a neutrino of energy $E$ into 3 particles of energy $E/3$ adds 3 particles to the spectrum which consists of $(E/(E/3))^{2}$ = 9 particles.
@@
For the aforementioned values of $\Lambda$ and $E$ one cannot neglect the electron mass values, so one should estimate the threshold effect, which is in general complicated for non-equal 3 particles in the final state. Consider for simplicity collinear kinematics, neutrino of energy and x-projection of momentum $(k+\frac{k^{3}}{2\Lambda^{2}},k)$ decays to a neutrino, an electron and a positron each of which takes $k/3$ momentum. Finally, the threshold of the effective decay reads,
$\Lambda_{thr}=\frac{E^{2}}{\sqrt{6}m_{e}}=3.9_{-2.7}^{+46.0}\times 10^{19}\,\mbox{GeV},\qquad n=2.$ (11)
For larger $\Lambda$ the process does not occur, for lower $\Lambda$ the calculation is similar to the neutrino splitting. Carmona et al [8] calculated the corresponding width in case of zero electron mass; the width is that of eq. (7) multiplied by $0.13$. Finally, for our values of parameters $\nu\to\nu e^{+}e^{-}$ is subleading with respect to neutrino splitting.
@@
Consider the opposite process, the production of muon neutrino. In the majority of models, neutrino is produced very fast in the process of pion decay, $\pi^{+}\to\mu^{+}\nu_{\mu}$. If the process is kinematically forbidden, no neutrino is produced. Two-particle decay threshold rule reads,
@@
From that we obtain a constraint,
$\Lambda=5.3_{-3.7}^{+5.8}\times 10^{17}\,\mbox{GeV},\qquad n=2.$ (13)
@@ Conclusion
We estimated constraints on LIV in the neutrino sector from the event KM3-230213A. It turns out that the best constraints came from the absence of the neutrino splitting. We considered separately the cases for neutrino of galactic, extragalactic, and cosmogenic origin. Since it is concluded in [2] that the galactic origin is unlikely, we may take the lower bound of the constraint interval (9) as the conservative $68\%$ CL bound, precisely
$\Lambda=\begin{cases}1.1\times 10^{30}\,\mbox{GeV},&n=1,\\ 1.1\times 10^{19}\,\mbox{GeV},&n=2.\end{cases}$ (14)
In order to obtain more precise statistically significant constraints, the models of galactic, extragalactic, or cosmogenic sources are required, as well as an accurate calculation of neutrino energy losses.
After the appearance of the first version of this manuscript, LIV analysis of the KM3-230213A event by the KM3NET collaboration appears [13]. This analysis is related to the scenario of constant superluminal neutrino velocity, or $n=0$ in our notations, see eq.(1). In this scenario, the neutrino splitting does not occur due to the energy-momentum conservation while the pair production $\nu\to\nu e^{+}e^{-}$ occurs over a certain threshold in the case of standard electrons.